OpenStudy (anonymous):

Temperature is usually measured in degrees Celsius or degrees Fahrenheit. In order to convert from degrees Fahrenheit to degrees Celsius, the following formula is used. °C = 5 9 (°F - 32) Which of the following answer choices correctly rearranges the above formula so that degrees Celsius can be converted to degrees Fahrenheit?

sammixboo (sammixboo):

What are your choices so I can help :P

OpenStudy (anonymous):

A: °F =5/9°C+32 B: °F=9/5(°C - 32) C: °F=9/5°C+32 D: °F=5/9(°C - 32)